Down Under Boomerang, Inc., is considering a new 3-year expansion project that requires an initial fixed asset investment of $2.31 million. The fixed asset will be depreciated straight-line to zero over its 3-year tax life, after which it will be worthless. The project is estimated to generate $1,725,000 in annual sales, with costs of $635,000. The tax rate is 23 percent and the required return is 12 percent. What is the project’s NPV? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,310,000/3)=$770,000/year

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(1,725,000-635,000)(1-0.23)+(0.23*770,000)

=1,016,400

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=1,016,400[1-(1.12)^-3]/0.12

=1,016,400*2.401831268

=2,441,221.30

NPV=Present value of inflows-Present value of outflows

=2,441,221.30-$2,310,000

=$131,221.30(Approx).